# Gaussian Eigen Models for Human Heads

### v2

@@ 2.1 NeRF-based avatars
Numerous methods [90, 41, 63, 81, 85, 57, 11, 78, 25, 51, 34, 16, 91] capitalize on the speed and quality of 3DGS. Qian et al. [57] attach 3D Gaussians to the FLAME [40] mesh surface and apply a deformation gradient similar to Zielonka et al. [88] to orient the Gaussians according to the local Frenet frames of the surface. This method, however, does not utilize any information about expressions and, thus, struggles with pose-dependent changes (e.g., wrinkles, self-shadows) and, despite high-quality results, retrieves only a global static appearance model. Li et al. [41] use a StyleUNet-like CNN architecture [74] to regress front and back Gaussian maps. Employing the powerful CNN network on position maps, they achieve impressive results for human bodies with effects like pose-dependent wrinkle formation.
@@ 3.1 Gaussian Eigen Model (GEM)
Given $\boldsymbol{D}=\{\boldsymbol{G}_{0},...,\boldsymbol{G}_{N-1}\}$, we build a personalized eigenbasis model, which is called GEM. We compute a statistical model for each Gaussian modality separately. Specifically, we create individual bases for rotation $\text{rotation }\boldsymbol{B}_{\theta},\text{position }\boldsymbol{B}_{\phi},\text{opacity }\boldsymbol{B}_{\alpha}$, and scale $\text{scale }\boldsymbol{B}_{\sigma}$ with respective means $\vec{\mu}_{\theta}$, $\vec{\mu}_{\phi}$, $\vec{\mu}_{\alpha}$ and $\vec{\mu}_{\sigma}$ via Principle Complonent Analysis (PCA) [26]. Note that the color $\boldsymbol{C}$ is optimized globally and, thus, acts as a classical texture without the need to apply PCA. A face model instance $\boldsymbol{G}$ is represented as a linear combination of these bases:
@@
We refine the base vectors for around 30k iterations. To ensure that the individual bases stay orthonormal throughout this refinement, every 1k steps, we orthogonalize the bases using QR decomposition. This refinement improves the training PSNR errors from 34.75dB to 36.68dB and 36.85dB for the training steps 0k, 5k, and 30k, respectively.
@@ 3.2 Image-based Animation
Expressions for a GEM are fully defined by their coefficients $\mathbf{k}_{\theta}$, $\mathbf{k}_{\phi}$, $\mathbf{k}_{\theta}$ and $\mathbf{k}_{\sigma}$. This is a similar idea to codec avatars [45], however, our approach does not need additional pixel shaders in the form of a small regression MLP. There are several ways to obtain the coefficients of a GEM, for example, one can employ analysis-by-synthesis-based optimization or regression. Analysis-by-synthesis [3] is the backbone of current avatar methods, as they use photometric or depth-based face trackers to sequentially optimize the coefficients of the underlying 3DMMs like FLAME [71, 88, 19] which is typically slow. As a fast, but more imprecise alternative, regressors like DECA [10] or EMOCA [7] can be used which are built on a ResNet backbone and regress FLAME parameters directly from an image.
As we aim for fast and expressive motion capturing, we build a personalized GEM regressor, that utilizes a pretrained EMOCA model. We apply several modifications to the EMOCA model, see 4. We use intermediate features of the pre-trained EMOCA network denoted as $\Theta(\mathbf{I}_{i})$ where $\mathbf{I}_{i}$ is the current image. EMOCA’s architecture comprises two ResNet networks; one to extract expression features $\mathbf{f}_{expr}\in\mathbb{R}^{2048}$ and the second for shape $\mathbf{f}_{shape}\in\mathbb{R}^{2048}$, both are followed by final MLPs to regress corresponding FLAME parameters. As we do not rely on FLAME, we remove the last hidden layer of the final MLP obtaining two feature vectors which we combine into one $\mathbf{f}\in\mathbb{R}^{2\times 1024}$ vector. For these features, we build a PCA layer with a basis denoted as $\mathbf{\hat{R}}$ using the training frames from five frontal cameras of NeRSemble. Note that we use relative features $\mathbf{r}=\mathbf{f}-\mathbf{f}_{neutral}$ in this PCA layer. The neutral reference frame $\mathbf{f}_{neutral}=\Theta(\mathbf{I}_{neutral})$ to compute these relative features is selected manually from the video, similar to Face2Face [71]. During training, for each frame, we project $\mathbf{r}$ onto the PCA manifold using the first $50$ principal components to restrict and regularize training. Finally, we use their corresponding PCA coefficients:
@@ 4 Results
We evaluate GEM on the NeRSemble [33], where tracked meshes [57] and synchronized images from 16 cameras with a resolution of $802\times 550$ are available. Our baselines are Gaussian Avatars (GA[57] which is neural network-free (Gaussians are attached to the FLAME model), our implementation of Animatable Gaussians (AG)  [41] which is based on CNN-predicting Gaussian maps, and INSTA [88] which uses dynamic NeRF [46]. For the comparison, we present both of our appearance models, the StyleUNet-based architecture (Ours Net) and the distilled linear Eigen model (Ours GEM) which we evaluate using analysis-by-synthesis fitting to the target images following [89, 71, 3]. Additionally, we present cross-reenactment results based on our coefficient regressor, compared to the baselines that use FLAME meshes regressed by DECA [10]. Relative expression transfer based on ground truth meshes [57] can be found in the supp. mat. All of the methods are evaluated using several image space metrics on novel expressions and novel views, following the test and novel-view split of Qian et al. [57]. For our GEM models, we use 50 components distilled from $256^{2}$ textures which give around 60k active Gaussians. Animatable Gaussians [41] uses a similar amount of primitives for front and back textures and Gaussian Avatars [57] around 100k Gaussians.
@@
Refer to captionSource ActorOurs GEMOurs NetGA [57]AG [41]INSTA [88]
@@ 4.3 GEM Ablation Studies
We are interested in the compression error introduced by the projection on different amounts of principal components used in GEM, also concerning the memory consumption. For instance, our smallest model weighs as little as 7MB using only 10 components of the eigenbasis. This is almost 12 times less than our smallest CNN-based model and almost 70 times less than Animatable Gaussians [41]. In contrast to neural networks, we can easily trade quality over size which is very useful in the context of different commodity devices with reduced compute capabilities. Table 4 presents how compression affects the quality of reconstruction, where we evaluate a sequence with $\sim 1$k frames for a single actor under a novel view. As expected, using only 10 components impacts the quality the most, however, the results are still of high quality, see Figure 8.
Gaussian Avatars [57] offers a small size of the stored Gaussians cloud, ranging from 5MB, and 14MB without the FLAME model for $128^{2}$, $256^{2}$ Gaussians, respectively. However, the quality of reconstruction lacks wrinkle details and sharpness as can be seen in Figure 12. In comparison to Gaussian Avatars [57], our model does not require FLAME during inference which is an additional 90MB.
@@ 5 Discussion
We design a universal method capable of distilling 3DGS-based avatar solutions into a lightweight representation, GEM, provided that normalized input across training frames is available. The only requirement to successfully distill GEM is to have a dataset with Gaussian-image pairs across the training sequences. Our results show that a compact representation of the linear basis produces state-of-the-art results in terms of quality and speed. Note that to achieve wrinkle-level details, the generator itself has to produce high-quality outputs. Our distillation technique can be applied to existing methods, making them lightweight and compact. GEM is well-suited for commodity devices, generating Gaussian primitives by a simple linear combination of the basis vectors. This potential has promising implications for tasks like holoportation, audio-driven avatars, and virtual reality.
The PCA-based GEM models have a global extent which is useful for some applications, but it also means that we cannot control local changes and produce more combinations of local features. Thus, further work could include incorporating a localized PCA basis [48] for better avatar control, which could potentially enable a wider range of expressions outside the training set. Another limitation is the personalization, for new subjects a new representation has to be learned from multi-view data. An interesting future avenue is to create a statistical model across subjects.
@@ 6 Conclusion
We have proposed Gaussian Eigen Models for Human Heads, a linear appearance model that represents photo-realistic head avatars. The simplicity of this appearance model results in massively reduced compute requirements in comparison to CNN-based avatar methods. Moreover, the lightweight representation could improve the management, sharing, and applicability of avatars. Our distillation approach can be applied to existing methods, making them available for compression. We demonstrate how GEMs can be used in scenarios like self-reenactment and cross-person animation, even in real-time.
@@ Acknowledgement
The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting WZ. JT is supported by the Google Gift Award.
@@ Appendix B Appearance Maps Generator
where $\boldsymbol{D}_{i}\in\mathbb{R}^{11\times H\times W}$ is a map containing the delta for positions $\Delta_{pos}\in\mathbb{R}^{3}$, rotation $\Delta_{rot}\in\mathbb{R}^{4}$, scale $\Delta_{scale}\in\mathbb{R}^{3}$, and opacity $\Delta_{op}\in\mathbb{R}$. The colors $\boldsymbol{C}\in\mathbb{R}^{3\times H\times W}$ of the Gaussians are predicted using the normal map $\phi(\boldsymbol{M})$ of the canonical mesh $\boldsymbol{M}$ ($\phi$ is the normal map extractor from a mesh). Similar to Animatable Gaussians [41], we use a PCA layer $\Gamma_{\text{PCA}}$ which serves as a low-pass regularization filter for the input. $\Gamma_{\text{PCA}}$ is built by using PCA on the meshes $\boldsymbol{M}_{i}$ for the training frames and 16 principle components are used as the basis. For the training, we use the projection of each incoming mesh $\boldsymbol{M}_{i}$ on the PCA manifold $\boldsymbol{M}_{PCA_{i}}=\Gamma_{\text{PCA}}(\boldsymbol{M}_{i})$.
@@ Appendix C Compression Ablation Study
Expanding upon this foundation, GEM proposes a novel technique involving an ensemble of eigenbases of 3D Gaussian attributes for achieving a photorealistic human head appearance. This representation exhibits significant adaptability concerning both quality and size, leveraging a fundamental trait of linear basis that proves beneficial when applied to diverse devices with varying capabilities in digital human applications. Figures 17 and 18 illustrate the qualitative and quantitative results of GEM. Notably, even under substantial compression (utilizing only ten principal components), our approach consistently yields high-quality outcomes. More examples can be found in Figures 19, 20, 21 and 22.
@@ Appendix D Human Head Avatar Compression
Figure 13: GEM applied on different avatar methods (AG and GA) and optimized using analysis-by-synthesis. Our method is universal and can be successfully used on both; point clouds as well as textures to distill a lightweight avatar.

### v3

@@ 2.1 NeRF-based avatars
Numerous methods [91, 41, 64, 82, 86, 58, 11, 79, 25, 52, 34, 16, 92] capitalize on the speed and quality of 3DGS. Qian et al. [58] attach 3D Gaussians to the FLAME [40] mesh surface and apply a deformation gradient similar to Zielonka et al. [89] to orient the Gaussians according to the local Frenet frames of the surface. This method, however, does not utilize any information about expressions and, thus, struggles with pose-dependent changes (e.g., wrinkles, self-shadows) and, despite high-quality results, retrieves only a global static appearance model. 3D Gaussian blendshapes [46] controls an avatar by linearly interpolating between optimized blendshapes using 3DMM expression coefficents. However, this method depends on underlying 3DMM whereas GEM is a mesh-free representation. Li et al. [41] use a StyleUNet-like CNN architecture [75] to regress front and back Gaussian maps. Employing the powerful CNN network on position maps, they achieve impressive results for human bodies with effects like pose-dependent wrinkle formation.
@@ 3.1 Gaussian Eigen Model (GEM)
Given $\boldsymbol{D}=\{\boldsymbol{G}_{0},...,\boldsymbol{G}_{N-1}\}$, we build a personalized eigenbasis model, which is called GEM. We compute a statistical model for each Gaussian modality separately. Specifically, we create individual bases for rotation $\text{rotation }\boldsymbol{B}_{\theta},\text{position }\boldsymbol{B}_{\phi},\text{opacity }\boldsymbol{B}_{\alpha}$, and scale $\text{scale }\boldsymbol{B}_{\sigma}$ with respective means $\vec{\mu}_{\theta}$, $\vec{\mu}_{\phi}$, $\vec{\mu}_{\alpha}$ and $\vec{\mu}_{\sigma}$ via Principle Complonent Analysis (PCA) [26]. Note that the color $\boldsymbol{C}$ is optimized globally and, thus, acts as a classical texture without the need to apply PCA. To accurately learn dynamically moving Gaussians, we fixed the color to prevent it from dominating the image representation, otherwise, Gaussians could change their semantic meaning (e.g., a Gaussian could represent the lip in one state, and the teeth in the other deformation state). Keeping the semantic meaning of specific Gaussians across deformation states is crucial for applying a PCA afterward.
A face model instance $\boldsymbol{G}$ is represented as a linear combination of these bases:
@@
We refine the base vectors for around 30k iterations. To ensure that the individual bases stay orthonormal throughout this refinement, every 1k steps, we orthogonalize the bases using QR decomposition. This refinement improves the training PSNR errors from 34.75dB to 36.68dB and 36.85dB for the training steps 0k, 5k, and 30k, respectively. Throughout our experiments, we did not encounter overfitting issues with this scheme. The reconstruction metrics on two randomly selected test sequences with refinement are: PSNR: 31.51, LPIPS: 0.091, SSIM: 0.936; and without: PSNR: 31.38, LPIPS: 0.094, SSIM: 0.933.
@@ 3.2 Image-based Animation
Expressions for a GEM are fully defined by their coefficients $\mathbf{k}_{\theta}$, $\mathbf{k}_{\phi}$, $\mathbf{k}_{\theta}$ and $\mathbf{k}_{\sigma}$. This is a similar idea to codec avatars [45], however, our approach does not need additional pixel shaders in the form of a small regression MLP. There are several ways to obtain the coefficients of a GEM, for example, one can employ analysis-by-synthesis-based optimization or regression. Analysis-by-synthesis [3] is the backbone of current avatar methods, as they use photometric or depth-based face trackers to sequentially optimize the coefficients of the underlying 3DMMs like FLAME [72, 89, 19] which is typically slow. As a fast, but more imprecise alternative, regressors like DECA [10] or EMOCA [7] can be used which are built on a ResNet backbone and regress FLAME parameters directly from an image. We apply several modifications to the EMOCA model, see 4. We use intermediate features of the pre-trained EMOCA network denoted as $\Theta(\mathbf{I}_{i})$ where $\mathbf{I}_{i}$ is the current image. EMOCA’s architecture comprises two ResNet networks; one to extract expression features $\mathbf{f}_{expr}\in\mathbb{R}^{2048}$ and the second for shape $\mathbf{f}_{shape}\in\mathbb{R}^{2048}$, both are followed by final MLPs to regress corresponding FLAME parameters. As we do not rely on FLAME, we remove the last hidden layer of the final MLP obtaining two feature vectors which we combine into one $\mathbf{f}\in\mathbb{R}^{2\times 1024}$ vector. For these features, we build a PCA layer with a basis denoted as $\mathbf{\hat{R}}$ using the training frames from five frontal cameras of NeRSemble. Note that we use relative features $\mathbf{r}=\mathbf{f}-\mathbf{f}_{neutral}$ in this PCA layer. The neutral reference frame $\mathbf{f}_{neutral}=\Theta(\mathbf{I}_{neutral})$ to compute these relative features is selected manually from the video, similar to Face2Face [72]. During training, for each frame, we project $\mathbf{r}$ onto the PCA manifold using the first $50$ principal components to restrict and regularize training. Finally, we use their corresponding PCA coefficients:
@@ 4 Results
We evaluate GEM on the NeRSemble [33], where tracked meshes [58] and synchronized images from 16 cameras with a resolution of $802\times 550$ are available. Our baselines are Gaussian Avatars (GA[58] which is neural network-free (Gaussians are attached to the FLAME model), our implementation of Animatable Gaussians (AG)  [41] which is based on CNN-predicting Gaussian maps, and INSTA [89] which uses dynamic NeRF [47]. Note that all baselines require at least two stages: (i) construct the avatar, and (ii) get the parameters to drive it. Most of them use offline tracking with additional objectives like hair reconstruction [58, 15], which does not work for real-time applications despite the avatar model’s rendering being real-time. Importantly, in our approach, we introduce a third step, i.e., the construction of the eigenbasis (GEM), which only introduces negligible computational costs ($\sim$ 1 min) in comparison to the avatar reconstruction itself. For the comparison, we present both of our appearance models, the StyleUNet-based architecture (Ours Net) and the distilled linear Eigen model (Ours GEM) which we evaluate using analysis-by-synthesis fitting to the target images following [90, 72, 3]. Additionally, we present cross-reenactment results based on our coefficient regressor, compared to the baselines that use FLAME meshes regressed by DECA [10]. Relative expression transfer based on ground truth meshes [58] can be found in the supp. mat. All of the methods are evaluated using several image space metrics on novel expressions and novel views, following the test and novel-view split of Qian et al. [58]. For our GEM models, we use 50 components distilled from $256^{2}$ textures which give around 60k active Gaussians. Animatable Gaussians [41] uses a similar amount of primitives for front and back textures and Gaussian Avatars [58] around 100k Gaussians.
@@ 4.1 Image Quality Evaluation
Refer to captionSource ActorOurs GEMOurs NetGA [58]AG [41]INSTA [89]
@@ 4.3 GEM Ablation Studies
We are interested in the compression error introduced by the projection on different amounts of principal components used in GEM, also concerning the memory consumption. Additionally, we show that GEM is not topology-dependent. Figure 9 demonstrates that our method is able to handle different topologies (subject wearing glasses), despite utilizing a fixed UV space.
Figure 9: Despite fixed topology and predefined texture size GEM faithfully represents facial attributes like glasses.
For instance, our smallest model weighs as little as 7MB using only 10 components of the eigenbasis. This is almost 12 times less than our smallest CNN-based model and almost 70 times less than Animatable Gaussians [41]. In contrast to neural networks, we can easily trade quality over size which is very useful in the context of different commodity devices with reduced compute capabilities. Table 4 presents how compression affects the quality of reconstruction, where we evaluate a sequence with $\sim 1$k frames for a single actor under a novel view. As expected, using only 10 components impacts the quality the most, however, the results are still of high quality, see Figure 8. Gaussian Avatars [58] offers a small size of the stored Gaussians cloud, ranging from 5MB, and 14MB without the FLAME model for $128^{2}$, $256^{2}$ Gaussians, respectively. However, the quality of reconstruction lacks wrinkle details and sharpness as can be seen in Figure 13. In comparison to Gaussian Avatars [58], our model does not require FLAME during inference which is an additional 90MB.
@@ 5 Discussion
We design a universal method capable of distilling 3DGS-based avatar solutions into a lightweight representation, GEM, provided that normalized input across training frames is available. The only requirement to successfully distill GEM is to have a dataset with Gaussian-image pairs across the training sequences. Our results show that a compact representation of the linear basis produces state-of-the-art results in terms of quality and speed. Note that to achieve wrinkle-level details, the generator itself has to produce high-quality outputs. Our distillation technique can be applied to existing methods like [92], making them lightweight and compact. GEM is well-suited for commodity devices, generating Gaussian primitives by a simple linear combination of the basis vectors. This potential has promising implications for tasks like holoportation, audio-driven avatars, and virtual reality.
Limitations
: The PCA-based GEM models have a global extent which is useful for some applications, but it also means that we cannot control local changes and produce more combinations of local features. Thus, further work could include incorporating a localized PCA basis [49] for better avatar control, which could potentially enable a wider range of expressions outside the training set. Other limitations are; side-view generalization which results in unstable expressions and personalization. For new subjects a new representation has to be learned from multi-view data. An interesting future avenue is to create a statistical model across subjects.
@@ 6 Conclusion
We have proposed Gaussian Eigen Models for Human Heads, a linear appearance model that represents photo-realistic head avatars. The simplicity of this appearance model results in massively reduced compute requirements in comparison to CNN-based avatar methods. Although the idea is simple, it offers many interesting downstream applications. The lightweight representation could improve the management, sharing, and applicability of avatars. Moreover, GEM simplifies the process of online avatar animation from RGB images and increases flexibility by balancing memory and quality trade-offs through additional control over the number of eigenbases. Our distillation approach can be applied to existing methods, making them available for compression. We demonstrate how GEMs can be used in scenarios like self-reenactment and cross-person animation, even in real-time.
@@ Acknowledgement
The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting WZ. JT was supported by the ERC Starting Grant LeMo (101162081). All the data were processed outside Google.
@@ Appendix B Appearance Maps Generator
where $\boldsymbol{D}_{i}\in\mathbb{R}^{11\times H\times W}$ is a map containing the delta for positions $\Delta_{pos}\in\mathbb{R}^{3}$, rotation $\Delta_{rot}\in\mathbb{R}^{4}$, scale $\Delta_{scale}\in\mathbb{R}^{3}$, and opacity $\Delta_{op}\in\mathbb{R}$. The colors $\boldsymbol{C}\in\mathbb{R}^{3\times H\times W}$ of the Gaussians are predicted using the normal map $\phi(\boldsymbol{M})$ of the canonical mesh $\boldsymbol{M}$ ($\phi$ is the normal map extractor from a mesh). Similar to Animatable Gaussians [41], we use a PCA layer $\Gamma_{\text{PCA}}$ which serves as a low-pass regularization filter for the input. $\Gamma_{\text{PCA}}$ is built by using PCA on the meshes $\boldsymbol{M}_{i}$ for the training frames and 16 principle components are used as the basis. For the training, we use the projection of each incoming mesh $\boldsymbol{M}_{i}$ on the PCA manifold $\boldsymbol{M}_{PCA_{i}}=\Gamma_{\text{PCA}}(\boldsymbol{M}_{i})$. Optionally we use a feature texture that is concatenated with rasterized normals for the $\boldsymbol{D}_{i}$ conditioning.
@@ Appendix C Compression Ablation Study
Expanding upon this foundation, GEM proposes a novel technique involving an ensemble of eigenbases of 3D Gaussian attributes for achieving a photorealistic human head appearance. This representation exhibits significant adaptability concerning both quality and size, leveraging a fundamental trait of linear basis that proves beneficial when applied to diverse devices with varying capabilities in digital human applications. Figures 18 and 19 illustrate the qualitative and quantitative results of GEM. Notably, even under substantial compression (utilizing only ten principal components), our approach consistently yields high-quality outcomes. More examples can be found in Figures 20, 21, 22 and 23.
@@ Appendix D Human Head Avatar Compression
Figure 14: GEM applied on different avatar methods (AG and GA) and optimized using analysis-by-synthesis. Our method is universal and can be successfully used on point clouds and textures to distill a lightweight avatar.